The pressure of gas is recorded as 1.0 x 105 N/m2 at a temperature of 0°C. The cylinder is heated further till the thermometer records 150°C. What is the pressure of the gas?
1
Expert's answer
2020-09-29T06:47:59-0400

QUESTION # 135355

The pressure of gas is recorded as 1.0 x 105 N/m2 at a temperature of 0°C. The cylinder is heated further till the thermometer records 150°C. What is the pressure of the gas?

ANSWER

P1VI/T1 =P2V2/T2

Volume remains constant therefore

P1/T1 =P2/T2

P1 =1.0× 105 N/M2

T1 = 0℃ = 0 + 273 =273K

P2 =?

T2 = 150 + 273 = 423K

1.0× 105 N/M2 ÷ 273K = P2 ÷ 423K


P2 = (1.0× 105 N/M2 × 423K) ÷ 273K

    =154945.0549 N/M2

    = 1.5×105 N/M2